OpenStudy (anonymous):

Write an equation of a line that passes through the point (2, 3) and is parallel to the line y =3x/2+5. (5 points) y =-2/3x y = -2/3x− 7 y = 3/2x y = 3/2x+ 7

OpenStudy (anonymous):

@annas can you please help me?

OpenStudy (anonymous):

@Loser66 can you please help?:)

OpenStudy (loser66):

The latex doesn't show. What is the given line's equation?

OpenStudy (anonymous):

sorry I'm confused?

OpenStudy (loser66):

y = .....x + 5? I don't see it

OpenStudy (anonymous):

y =-2/3x y = -2/3x− 7 y = 3/2x y = 3/2x+ 7

OpenStudy (loser66):

Please, don't copy and paste. It doesn't work on my computer

OpenStudy (anonymous):

oh sorry i didn't know!

OpenStudy (anonymous):

y= -2/3x y= -2/3x -7 y= 3/2x y= 3/2x + 7

OpenStudy (anonymous):

@Loser66

OpenStudy (loser66):

Want me guide or my answer?

OpenStudy (anonymous):

both please!:)

OpenStudy (loser66):

the given equation is \(y = \dfrac{3}{2}x +5\) right?

OpenStudy (anonymous):

yep

OpenStudy (loser66):

the slope of it is 3/2, right?

OpenStudy (anonymous):

yes

OpenStudy (loser66):

hence, the required line has the same slope with it (3/2) , now you can pick either c or d.

OpenStudy (anonymous):

c?

OpenStudy (loser66):

Suppose c is the correct answer, then you need put the point in, if it satisfies the equation, then c is the right one, if it doesn't, we check d

OpenStudy (anonymous):

okay

OpenStudy (loser66):

to c, \(y = \dfrac{3}{2}x\), I replace the point (2,3) that is y = 3, x =2 so \(3 = \dfrac{3}{\cancel 2}\cancel 2\\3=3\) Bingo, hence c is the correct one

OpenStudy (loser66):

Then, no need to check d. C is the correct answer

OpenStudy (anonymous):

thank you so very much!